OpenStudy (jannine24):

1. Mrs. Lewis is returning back to the workforce after three years of raising her son Joey. Before she goes back to work she must find suitable daycare for Joey, one that will not break the bank either. After months of looking she has narrowed the field to two final selections Kiddie Castle which is a daycare facility about half way between home and work, the other is Cultural Au Pair in which the person comes to your home and takes care of the child there. Kiddie Castle charges a weekly rate of $112 plus $5.00 per hour of care for Joey. Cultural Au Pair charges a hourly rate of $7.00, you can

OpenStudy (anonymous):

satellite help me?

OpenStudy (jannine24):

who

OpenStudy (jannine24):

i have the 2 equations Kiddie Castle: Y = $112 + $5.00(x) Cultural Au Pair: Y = $7.00(x) y is the total cost and x number of horus

OpenStudy (anonymous):

What question are you asking ? Are you asking which one would be better (financially) ? It would depend on how many hours a week she works.

OpenStudy (jannine24):

i need to find out hwne these two would be the same cost

OpenStudy (anonymous):

112 + 5x = 7x 112 = 7x - 5x 112 = 2x 56 = x 112 + 5(56) = 7(56) 112 + 280 = 392 392 = 392 (correct) The only way they would be the same cost is if she worked 56 hrs per week Is this what you are asking ?

OpenStudy (jannine24):

yes that is what i need to find out

OpenStudy (jannine24):

so i subsitute 7 per hour into the first problem for y then solve

OpenStudy (anonymous):

no....you take the equations that were made from the problem..and set them equal to each other, then solve for x (hours worked). For them to be equal, x would be 56 hrs.

OpenStudy (anonymous):

you substitute 56 in for x to find y.

OpenStudy (jannine24):

i see now ty for your help

OpenStudy (anonymous):

your welcome :)